# Causal Attribution via Activation Patching

### 13 Mar 2026 v1

@@ 4.3 Results
4.3.1 Quantitative
Faithfulness.
Table 1 reports faithfulness metrics across four ViT models. Table 3 further presents the same faithfulness metrics on ImageNet-Hard [36] and Oxford-IIIT-Pet [24]. Our method consistently achieves the lowest Deletion AUC and the highest Insertion AUC on most backbones. In particular, our approach improves the Insertion-Deletion with a notable margin compared to prior methods, indicating that the regions ranked as important by our attribution maps more accurately reflect the model’s true decision-making process.
@@ 4.3.1 Quantitative
Localization results are reported in Table 2. Our method achieves the best overall performance across most models. In particular, the higher PG and AUPR1 values indicate that the regions identified as most relevant are consistently located within the annotated foreground areas. Moreover, strong AUPR0 scores demonstrate effective suppression of background activations. These results confirm that our approach produces spatially precise attribution maps that better align with ground-truth object regions. Further quantitative results are provided in Appendix A.1.
Figure 2: Qualitative attribution comparison between different methods for a representative ImageNet sample. CAAP produces more compact and well-localized attributions for the target class than the baselines. More examples are provided in the Appendix A.2.
Figure 3: Qualitative attribution comparison in a representative image containing several objects using CLIP-L/14. CAAP produces compact and class-specific attributions aligned with the queried class. More examples are provided in the Appendix A.2.
4.3.2 Qualitative
Qualitative results are provided in Figs 2 and 3. As illustrated, baseline methods often produce diffuse or fragmented attribution maps that extend beyond the target object into the background, or highlight only a small subset of discriminative patches. In contrast, our method generates more compact, coherent, and semantically meaningful attributions that align closely with the target object. Specifically, our attributions better capture object boundaries and maintain consistency across patches, resulting in fewer spurious activations on irrelevant background regions. The improved localization quality observed in the visualizations is consistent with our superior AUPR1, AUPR0 and PG scores reported in Table 2. Additional qualitative results are provided in Appendix A.2.
@@ 4.4 Ablation Studies
We ablate three design choices in CAAP on ImageNet across all four ViT backbones: (i) the type of the target’s blank patches, (ii) the type of the selection operator, and (iii) the choice of intervened layers. Additional analysis regarding efficient implementation via precomputation of blank-context attention, comparison with input-based approaches and ablation analysis on convolutional architectures are provided in Appendices B.1B.2 and B.3.
@@
We first ablate the type of blank target image patches used when constructing the patched target, varying the blank content among Black, Blur Noisy, Mean, Noisy, and White. Here, Mean fills the blanks with the dataset per-channel mean, Noisy fills them with i.i.d. Gaussian noise, and Blur Noisy first adds i.i.d. Gaussian noise to the target and then applies a blur filter on it. These choices span both in-distribution and out-of-distribution targets, since Mean matches dataset statistics while Black, White, and the noise-based variants introduce increasingly OOD blank content. Fig. 4 reports results on ImageNet across all four ViT backbones. Faithfulness is largely insensitive to this choice, as Insertion, Deletion, and Insertion-Deletion remain close across blank patch types for each model. AUPR0 and AUPR1 vary only marginally, while PG shows modest differences across backbones. We use White blanks as the default setting in all experiments.
@@
We next ablate the spatial support of the intervention, motivated by the fact that ViT patch tokens absorb information from nearby patches via self-attention. Fig. 5 compares four neighborhood variants. Overall, the no-padding variant performs the worst, while all padded alternatives consistently improve both faithfulness and localization across different architectures. This result aligns with our selection operator design (Section 3.1), confirming that because features spread spatially, patching a single token often misses context, and spatial padding improves attribution quality. We use the Radius-1 Box setting, which includes a $3\times 3$ patch region, as the default in our experiments.
@@
To analyze the effect of layer depth, we first examine attention patterns that govern information flow (Section 3.2). Fig. 6 shows the mean intra- and extra-object attention weights across layers, indicating that $\mathbb{E}_{i}[\mathbb{E}_{j\in\mathcal{O}}[\alpha_{ij}^{l}]-\mathbb{E}_{j\notin\mathcal{O}}[\alpha_{ij}^{l}]]$ peaks in the middle layers. To identify the optimal patching depth, we execute a cumulative layer sweep (Fig. 7), progressively patching all transformer blocks from the first up to a chosen cutoff layer. Across all ViT backbones, faithfulness consistently peaks at intermediate layers; interventions on early blocks mainly inject low-level features with limited class evidence, while extending patching into later blocks reduces reliability, as late representations are dominated by global mixing and weaker spatial specificity. Based on this observation, which aligns with the analysis in Section 3.2, we fix $l_{e}=\lceil\tfrac{2}{3}L\rceil$ as the default setting for all architectures, as it balances intra-object grouping and extra-object contextualization and is near-optimal across models.
5 Related Work
We review prior work on explaining Vision Transformers (ViTs) through three commonly used sources of explanatory signal: gradients, attention, and interventions based on perturbations.
@@ 5 Related Work
[…]
@@
Because ViTs rely on self-attention, many explanations derive saliency from attention weights and their compositions. Attention Flow and related rollout-style approaches compose attention across layers to propagate token influence [1]. However, attention weights alone do not necessarily align with causal importance [16, 17], motivating approaches that incorporate class information. Transformer Attribution propagates relevance through attention blocks using gradient-informed updates [7], and AttCAT combines features, gradients, and attention to produce token-level explanations [26]. Transition Attention Maps (TAM) models information flow through attention as a Markov process [41], while Grad-SAM leverages gradient self-attention maps to highlight influential inputs within attention units [2]. Graph-structured attention explanations also exist; MUTEX constructs a multiplex network over attention patterns across layers to generate patch heatmaps [18].
@@
Perturbation approaches estimate importance by modifying inputs or tokens and observing output changes. RISE aggregates responses over randomized masks [25], and optimization-based masking seeks minimally sufficient regions [12, 11]. For ViTs, ViT-CX builds causal explanations by masking patch embeddings and aggregating their effects on predictions [40], and Transformer Input Sampling (TIS) perturbs by sampling or dropping input tokens to avoid introducing replacement artifacts [10]. AtMan probes relevance by manipulating attention interactions efficiently [8]. Metric-Driven Attributions (MDA) directly optimizes attributions to score well under faithfulness metrics such as insertion and deletion [38].
Causal interventions and activation patching.
Causal viewpoints formalize attribution as the effect of interventions rather than local sensitivity [6]. In mechanistic interpretability, activation patching (interchange interventions or causal tracing) replaces internal activations from a source run into a target run while keeping the remaining computation fixed, enabling direct tests of which internal representations carry information used for a prediction [42].
6 Discussion
Empirically, CAAP improves faithfulness across multiple backbones, including CLIP, DINOv2, ViT, and DeiT3, by achieving lower Deletion AUC and higher Insertion AUC. Localization metrics such as AUPR1, AUPR0, and Pointing Game also improve, indicating that patch-level interventions better align with object regions rather than highlighting only a few highly discriminative patches. These results suggest that activation patching is not only a well-founded causal probe, but also an effective and scalable tool for identifying spatial evidence in ViTs. However, this method also has limitations, as it is largely fitted to transformer architectures. The activation patching reflects marginal patch effects and does not completely model higher-order interactions between patches. Extending the approach beyond this setting, including to other architectures or tasks such as multimodal or generative models, remains an open direction.
Appendix A Broader Experiments
In this section, we present additional experimental results that further expand the empirical evaluation in the main paper. In particular, we extend the study to other model scales, including base and, where available, small variants, as well as the largest variants of the considered architectures (huge/giant models). Additional quantitative results are reported in Section A.1, and further qualitative examples are provided in Section A.2. We also provide representative insertion and deletion curves for the large variants of the backbones in Figs 8 and 9.
@@ A.1 Quantitative Results
A.2 Qualitative Results
@@
Appendix B Broader Ablation Analysis
B.1 Efficient implementation of CAAP
The efficient implementation reduces the cost of computing attention between each classification token ($\mathrm{CLS}$) and the blank-context tokens. For each head, consider the self-attention update of a $\mathrm{CLS}$ token at layer $l$,
@@ B.1 Efficient implementation of CAAP
In the patched context, tokens are divided into source-patched tokens $j\in\mathcal{S}(p_{i})$ and blank tokens $j\notin\mathcal{S}(p_{i})$. For the blank tokens we observe that the dot products $\langle q_{\mathrm{CLS}}^{c^{*}(p_{i})},k_{0,j}^{l}\rangle$ are nearly identical to those obtained in the blank context $c^{\prime}$. Therefore we precompute two quantities for every layer and attention head.
First, we compute the normalization contribution of blank tokens,
@@
which is reused as a fixed term in the softmax denominator. Second, we compute a prototype value vector for the blank tokens,
@@
where $\mathcal{B}$ denotes the set of blank tokens. During attribution we reuse $Z_{0}^{l}$ and $\bar{v}_{0}^{l}$ and compute only the terms involving the source-patched tokens dynamically. This avoids repeated evaluation of dot products and value aggregations for blank tokens across images.
This efficient implementation substantially reduces the computational complexity. In a naive implementation, each attribution requires patching operations and attention recomputation over all tokens, leading to approximately $\mathcal{O}(N^{3})$ operations for an image with $N$ tokens ($N$ patch interventions combined with $N\times N$ attention interactions). With the proposed implementation, the attention contributions of blank tokens are precomputed and reused as a constant term ($\mathcal{O}(1)$) in the attention computation involving the $\mathrm{CLS}$ token and the patched tokens. As a result, only the $N$ patch interventions are evaluated dynamically, reducing the overall computational cost to approximately $\mathcal{O}(N)$. Tables 10 and 11 compare the original CAAP implementation with the efficient version (E-CAAP). The results show that both variants produce nearly identical performance across all four ViT backbones. Faithfulness metrics (Deletion, Insertion, and Ins-Del) remain very close for all models, and the localization metrics (AUPR1, AUPR0, and PG) show the same pattern. These results indicate that the approximation used for the blank-token attention does not noticeably affect attribution quality while substantially reducing the computational cost.
@@
B.2 Input-Based Intervention Baselines
Table 12: Faithfulness quantitative results for Simple Insertion, Simple Deletion, and CAAP on ImageNet. Deletion ($\downarrow$), Insertion ($\uparrow$), and Ins-Del ($\uparrow$) are reported. Best is bold, second best is underlined in each column.

### 17 May 2026 v2

@@ 4.3 Quantitative Results
Localization results are reported in Table 2. CAAP achieves the best overall performance across most backbones. In particular, the higher PG and AUPR1 values indicate that the regions identified as most relevant are consistently located within the annotated foreground areas. Moreover, strong AUPR0 scores demonstrate effective suppression of background activations. These results confirm that CAAP produces spatially precise attribution maps that better align with ground-truth object regions. Further quantitative results are provided in Appendix B.1, and evaluation with compactness metrics is reported in Appendix B.2.
Figure 2: Qualitative attribution comparison between different methods for a representative ImageNet sample. CAAP produces more compact and well-localized attributions for the target class than the baselines. More single-object qualitative examples are provided in the Appendix B.3 (Fig. 13– 42).
Figure 3: Qualitative attribution comparison between different methods for an image containing multiple objects. CAAP produces compact and class-specific attributions aligned with the queried object class. More multi-object qualitative examples are provided in Appendix B.3 (Fig. 4352).
4.4 Qualitative Results
Qualitative results are provided in Figs. 2 and 3. As illustrated, baseline methods often produce diffuse or fragmented attribution maps that extend beyond the target object into the background, or highlight only a small subset of discriminative patches. In contrast, CAAP tends to generate more compact, coherent, and semantically meaningful attributions that align closely with the target object. Specifically, CAAP’s attributions better capture object boundaries and maintain consistency across patches, resulting in fewer activations on irrelevant background regions. The improved localization quality observed in the visualizations is consistent with our superior AUPR1, AUPR0 and PG scores reported in Table 2. Additional qualitative results are provided in Appendix B.3.
@@ 4.5 Ablation Studies
We ablate three design choices in CAAP on ImageNet: (i) the type of the target’s blank patches, (ii) the type of the selection operator, and (iii) the choice of intervened layers. Results are shown for CLIP-L/14, while results on other ViT backbones (ViT-L/16, DINOv2-L/14, DeiT3-L/16) are provided in Appendix D.1. Additional analysis comparing CAAP with input-based approaches for both ViT and convolutional architectures is provided in Appendices D.2 and D.3. These results further demonstrate the effectiveness of CAAP on convolutional architectures and the superiority of activation patching over input-based interventions.
@@
We first ablate the type of blank target image patches used when constructing the patched target, varying the blank content among Black, White, Mean, Noisy, and Blur Noisy. Here, Black and White use all-black and all-white blank patches, respectively, Mean fills the blanks with the dataset per-channel mean, Noisy fills them with i.i.d. Gaussian noise, and Blur Noisy first adds i.i.d. Gaussian noise to the target and then applies a blur filter. Fig. 4(a) shows that faithfulness is largely insensitive to this choice, as Insertion, Deletion, and Insertion-Deletion remain close across blank types. AUPR0 and AUPR1 vary only marginally, while PG shows modest differences. We use White blanks as the default setting in all experiments.
@@
We next ablate the spatial support of the intervention, motivated by the fact that ViT patch tokens absorb information from nearby patches via self-attention. Fig. 4(b) compares four neighborhood variants. The no-padding variant performs the worst, while all padded alternatives improve both faithfulness and localization. This aligns with our selection operator design (Section 3.1), indicating that patching only a single token into a large number of blank patches often does not provide sufficient signal, while spatial padding strengthens the informative signal and improves attribution quality. We use the Radius-1 Box setting, which corresponds to a $3\times 3$ region, as the default in our experiments.
@@
To analyze the effect of layer depth, we first examine attention patterns that govern information flow (Section 3.2). Fig. 5(a) shows that $\mathbb{E}_{i}[\mathbb{E}_{j\in\mathcal{O}}[\alpha_{ij}^{l}]-\mathbb{E}_{j\notin\mathcal{O}}[\alpha_{ij}^{l}]]$ peaks in the middle layers. To identify the optimal patching depth, we execute a cumulative layer sweep (Fig. 5(b)), progressively patching all transformer blocks from the first up to a cutoff layer. Insertion AUC shows less sensitivity to the intervention depth than Deletion AUC. Overall, faithfulness peaks at intermediate layers: interventions on early blocks mainly inject low-level features with limited class evidence, while extending patching into later blocks reduces reliability, as late representations are dominated by global mixing and weaker spatial specificity. Based on this observation, which aligns with the analysis in Section 3.2, we fix $l_{e}=\lceil\tfrac{2}{3}L\rceil$ as the default setting for all architectures. This choice balances intra-object grouping and extra-object contextualization and performs near-optimally across models. Importantly, the performance (particularly Insertion AUC) remains stable across a broad range of intermediate layers, indicating that the method is not highly sensitive to the exact cutoff choice (results on other backbones are provided in Appendix D.1).
(a) Target blank type.
(b) Selection operator (padding).
Figure 4: Ablation studies of activation patching context characteristics on ImageNet using CLIP-L/14. (a) Effect of target blank type (Black, Blur Noisy, Mean, Noisy, White). (b) Effect of spatial support in the selection operator (No Padding, Radius-1 Box, Radius-2 Box, Radius-1 Manhattan). Faithfulness (Insertion, Deletion, Ins-Del) and localization (PG, AUPR1, AUPR0) are reported. Results on other ViT backbones are provided in Appendix D.1.
(a) Attention statistics across layers.
(b) Layer-wise intervention sweep.
Figure 5: Layer analysis on ImageNet using CLIP-L/14. (a) Mean intra-object, inter-object, and object-to-background attention across layers. In early layers, intra-object attention increases and dominates, indicating strong object-level grouping. Around the middle layers, this gap reaches its peak. In later layers, intra-object attention decreases while object-to-background attention increases, reflecting growing extra-object contextualization and reduced spatial specificity. (b) Cumulative intervention depth sweep showing faithfulness metrics. Intermediate layers provide the best trade-off between object-level grouping and global contextual mixing.
5 Discussion and Limitations
Empirically, CAAP improves faithfulness across multiple backbones, including CLIP, DINOv2, ViT, and DeiT3, by achieving lower Deletion AUC and higher Insertion AUC. Localization metrics such as AUPR1, AUPR0, and Pointing Game also improve, indicating that patch-level interventions better align with object regions rather than highlighting only a few highly discriminative patches. These results suggest that activation patching is not only a well-founded causal probe, but also an effective and scalable tool for identifying spatial evidence in ViTs. However, this method also has limitations, as it is largely fitted to transformer architectures. Although activation patching preserves higher-order interactions between patches from the source context and avoids OOD behavior in its internals, the patched context does not follow the standard activation distribution. While this does not degrade model performance in practice, and ViTs appear robust to this kind of OOD activation, it should still be considered in mechanistic interpretability studies. Extending the approach beyond this setting, specifically to multimodal decoder architectures or generative models, remains an open direction.
Appendix A Related Work
We review prior work on explaining Vision Transformers (ViTs) through three commonly used sources of explanatory signal: gradients, attention, and input-level perturbations.
@@ Appendix A Related Work
Gradient-based approaches estimate relevance using local sensitivity of the class score with respect to inputs or intermediate features. Representative methods include activation-difference propagation and path-integrated gradients, such as DeepLIFT and Integrated Gradients [32, 36], with noise-reduction variants such as SmoothGrad [34]. CAM-style localization derives class-discriminative maps from gradient-weighted or score-weighted features, including Grad-CAM, Grad-CAM++, and Score-CAM [31, 5, 40]. For ViTs, LibraGrad shows that nonlinear layers can make gradient-based attribution unfaithful, and proposes post-hoc gradient corrections that approximately mitigate this issue for common nonlinear operations in Transformers [21]. Layer-aggregation variants such as SkipPLUS further strengthen explanations by reducing the effect of noisy early layers when aggregating gradient-based signals [22]. In contrast to gradient- and relevance-propagation methods, which trace local sensitivity or backward relevance through the computation graph, CAAP directly intervenes on the forward activations of contextualized patch tokens.
@@
Because ViTs rely on self-attention, many explanations derive saliency from attention weights and their compositions. Attention Flow and related rollout-style approaches compose attention across layers to propagate token influence [1]. However, attention weights alone do not necessarily align with causal importance [17, 18], motivating approaches that incorporate class information or additional model signals. Transformer Attribution propagates relevance through attention blocks using gradient-informed updates [7], and AttCAT combines features, gradients, and attention to produce token-level explanations [27]. Transition Attention Maps (TAM) models information flow through attention as a Markov process [42], while Grad-SAM leverages gradient self-attention maps to highlight influential inputs within attention units [2]. Graph-structured attention explanations also exist; MUTEX constructs a multiplex network over attention patterns across layers to generate patch heatmaps [19]. AtMan manipulates attention mechanisms in Transformers to produce relevance maps without relying on backpropagation, using a memory-efficient token-search procedure based on neighborhoods in embedding space [8]. Unlike attention-based methods, which use token-mixing patterns or attention manipulations as explanatory signals, CAAP measures the effect of patch-associated representations on the model output through an explicit activation intervention.
@@
Perturbation approaches estimate importance by modifying inputs, masks, or early token representations and observing output changes. RISE aggregates responses over randomized input masks [26], and optimization-based masking seeks minimally sufficient or necessary regions [12, 11]. For ViTs, ViT-CX builds causal explanations from patch embeddings and their causal impact on the model output [41], and Transformer Input Sampling (TIS) perturbs the model by sampling input tokens, using the variable-token property of Transformers to avoid introducing replacement artifacts [10]. Metric-Driven Attributions (MDA) directly optimizes attribution maps to score well under faithfulness metrics such as insertion and deletion [39]. These methods provide important perturbation-based or metric-driven views of ViT attribution, but their interventions are applied to external inputs, early token representations, or the attribution map itself. In contrast, CAAP intervenes on contextualized residual activations inside the ViT.
Positioning of CAAP.
CAAP can be viewed as an interventional attribution approach that assigns relevance by replacing internal activations and measuring the resulting target-class score. It complements masking- and sampling-based perturbation methods by moving the intervention from the input or token level to the representation level, and it complements attention-derived explanations by treating attention as part of the computation that forms the intervened states rather than as the explanatory signal itself. It is also distinct from metric-driven approaches that optimize attribution maps to fit evaluation criteria: CAAP derives each score from an explicit intervention within the classifier itself, without learning an external explainer, mask generator, or metric-specific surrogate. The key distinction is that CAAP directly tests the predictive effect of intermediate residual representations associated with image patches. Methods such as RISE [26] estimate importance by aggregating model responses over input-space masks, while ViT-CX [41] constructs explanations from patch embeddings and output-level causal scores. These methods therefore do not directly measure how the contextualized residual activations formed inside the transformer affect the final prediction. CAAP addresses this gap by inserting source-image residual activations into a controlled target computation and asking whether those internal patch representations support the target-class prediction. Thus, CAAP shifts the intervention site for ViT attribution from external or early model variables to internal representations that are closer to the model’s final decision computation.
Appendix B Broader Experiments
In this section, we present additional experimental results that further expand the empirical evaluation in the main paper. In particular, we extend the study to other model scales, including base and, where available, small variants, as well as the largest variants of the considered architectures (huge/giant models). Additional quantitative results are reported in Section B.1, and further qualitative examples are provided in Section B.3. We also provide representative insertion and deletion curves for the large variants of the backbones in Figs 6 and 7.
@@ B.1 Additional Quantitative Results
B.2 Compactness Metrics
In addition to faithfulness and localization, we evaluate attribution maps using compactness metrics that measure how concentrated the attribution mass is for single-object classification. Given nonnegative patch scores $a_{i}$ and normalized scores $p_{i}=a_{i}/\sum_{j}a_{j}$, we report normalized Entropy,
$H(p)=-\frac{1}{\log N}\sum_{i=1}^{N}p_{i}\log p_{i},$
and the Gini index,
$G=\frac{\sum_{i=1}^{N}(2i-N-1)a_{(i)}}{N\sum_{i=1}^{N}a_{i}},$
where $a_{(i)}$ are scores sorted in increasing order. These metrics quantify whether attribution is focused on a compact set of relevant regions rather than being diffuse. Table 10 presents the results across multiple ViT backbones. CAAP generally achieves lower entropy and higher Gini scores, indicating more compact and coherent attribution maps.
Table 10: Compactness quantitative results on ImageNet for ViT-L/16 [35], CLIP-L/14 [28], DINOv2-L/14 [23], and DeiT3-L/16 [38]. Entropy ($\downarrow$) and Gini ($\uparrow$) are reported. Best is bold, second best is underlined in each column.
ViT-L/16 [35] CLIP-L/14 [28] DINOv2-L/14 [23] DeiT3-L/16 [38]
[…]
RISE [26] 97.93±0.04 24.58±0.24 98.03±0.03 24.80±0.18 98.35±0.03 22.65±0.18 98.02±0.04 24.02±0.23
Grad-CAM [31] 93.33±0.20 42.73±0.59 73.09±0.50 79.85±0.50 94.39±0.14 41.03±0.49 85.90±0.30 60.21±0.59
Rollout [1] 96.42±0.03 33.53±0.15 96.59±0.05 33.09±0.22 83.98±0.12 60.53±0.20 93.28±0.08 41.32±0.27
T-Attr [7] 88.93±0.12 53.99±0.30 89.15±0.41 55.24±0.88 95.89±0.12 32.75±0.37 81.01±0.31 69.68±0.42
ViT-CX [41] 97.14±0.07 27.79±0.44 97.63±0.07 25.87±0.43 97.31±0.05 28.51±0.32 95.90±0.10 33.95±0.44
TAM [42] 94.50±0.10 40.66±0.33 94.42±0.13 41.18±0.42 96.60±0.06 32.98±0.28 86.52±0.24 60.22±0.48
IxG+ [22] 99.42±0.02 10.31±0.22 99.55±0.02 8.85±0.19 99.47±0.01 10.74±0.17 96.51±0.20 18.93±0.57
Libra IxG+ [21] 97.57±0.09 23.50±0.46 98.99±0.05 12.77±0.35 98.52±0.06 19.10±0.40 94.88±0.17 34.49±0.60
MDA [39] 86.89±0.29 58.24±0.57 85.47±0.31 62.69±0.57 86.37±0.29 59.97±0.53 90.40±0.20 51.83±0.52
MUTEX [19] 98.40±0.03 21.63±0.18 98.38±0.02 22.48±0.18 97.58±0.05 27.08±0.31 98.40±0.02 21.55±0.17
[…]
B.3 Additional Qualitative Results
@@
Appendix C Efficient Implementation via Blank-Context Precomputation
A large fraction of the computation in CAAP arises from attention between each classification token ($\mathrm{CLS}$) and patch tokens originating from the blank image $x_{0}$. Let $K_{0}^{l}$ and $V_{0}^{l}$ denote the key and value matrices at layer $l$ computed from $x_{0}$. These quantities are fixed and do not depend on the source image $x$ or the selected patch $p_{i}$.
In the patched context $c^{*}$, the update of each classification token $\mathrm{CLS}_{i}$ follows the standard self-attention operation,
@@ Appendix C Efficient Implementation via Blank-Context Precomputation
Here, tokens are divided into source-patched tokens $j\in\mathcal{S}(p_{i})$ and blank tokens $j\notin\mathcal{S}(p_{i})$, with keys and values drawn accordingly.
Empirically, we observe that for blank tokens the dot products $\langle q_{\mathrm{CLS}}^{c^{*}(p_{i})},k_{0,j}^{l}\rangle$ are nearly identical to those obtained in the blank context $c^{\prime}$. As a result, the attention distribution over blank tokens is effectively invariant to the inserted source activations.
We therefore precompute two quantities for each layer and attention head using the blank context. First, the normalization contribution of blank tokens,
@@
which is reused as a fixed term in the softmax denominator. Second, a prototype value vector for the blank tokens,
@@
where $\mathcal{B}$ denotes the set of blank tokens.
During attribution, we reuse $Z_{0}^{l}$ and $\bar{v}_{0}^{l}$ and compute only the terms involving the source-patched tokens dynamically. This avoids repeated evaluation of dot products and value aggregations for blank tokens across images.
This approximation substantially reduces computational complexity. A naive implementation requires approximately $\mathcal{O}(N^{3})$ operations for an image with $N$ tokens, due to $N$ patch interventions combined with $N\times N$ attention interactions. With precomputation, the contribution of blank tokens becomes a constant term ($\mathcal{O}(1)$), and only the $N$ patch interventions are evaluated dynamically, reducing the overall cost to approximately $\mathcal{O}(N)$.
Tables 11 and 12 compare the original CAAP implementation with the efficient version (E-CAAP). Both variants produce nearly identical results across all four ViT backbones. Faithfulness metrics (Deletion, Insertion, and Ins-Del) and localization metrics (AUPR1, AUPR0, and PG) remain very close, indicating that the approximation does not affect attribution quality while substantially reducing computational cost.
@@
Appendix D Broader Ablation Analysis
D.1 Ablation Analysis across ViT Models
Here we extend the reported ablation results to all four studied ViT backbones (CLIP-L/14, ViT-L/16, DINOv2-L/14, and DeiT3-L/16). Fig. 9 shows that the choice of blank target type has only a minor effect on both faithfulness and localization metrics across architectures, with White blanks performing consistently well. Fig. 10 demonstrates that adding spatial padding to the selection operator consistently improves attribution quality compared to patching a single token, supporting the use of local neighborhoods during activation patching. Fig. 11 analyzes attention statistics across layers and shows that intra-object attention dominates in the intermediate layers before later transitioning toward extra-object contextualization. Finally, Fig. 12 presents cumulative layer-wise intervention sweeps across all backbones, showing that faithfulness peaks at intermediate depths, validating our default choice of intervening up to approximately the first two-thirds of transformer layers.
D.2 Input-Based Intervention Baselines
Table 13: Faithfulness quantitative results for Input Insertion, Input Deletion, and CAAP on ImageNet. Deletion ($\downarrow$), Insertion ($\uparrow$), and Ins-Del ($\uparrow$) are reported. Best is bold, second best is underlined in each column.